X^3-2y^3+xy(2x-y)+y(x-y)+1=0


1
Expert's answer
2021-11-08T06:11:38-0500

Incomplete question:

No information is given that what needs to done here.

Let us solve for the derivative w.r.t "x" :

Let "f(x,y)=x^3-2y^3+xy(2x-y)+y(x-y)+1\\"

"=x^3-2y^3+2x^2y-xy^2+xy-y^2+1"

"f_x=\frac{\partial}{dx}[x^3-2y^3+2x^2y-xy^2+xy-y^2+1] \\
=3x^2+4xy-y^2+y"